OpenStudy (anonymous):

If 4/5+(3/10) = x+1 1/2 then x=

OpenStudy (anonymous):

If 4/5+(3/10) = x+1 1/2 then x=

OpenStudy (anonymous):

Start by putting all your fractions over a common denominator.

OpenStudy (anonymous):

\[\frac{4}{5}+\frac{3}{10}=x+1\tfrac{1}{2}\]? \[\frac{8}{10}+\frac{3}{10}=\frac{13}{10}=x+\frac{3}{2}\] \[\frac{13}{10}-\frac{15}{10}=x\] \[-\frac{2}{10}=-\frac{1}{2}=x\]